# Zero-shot Logical Query Reasoning on any Knowledge Graph

### 10 Apr 2024 v1

@@ 6 Discussion and Future Work
Limitations. The only parameterized component of UltraQuery is the projection operator and, therefore, limitations and improvement opportunities stem from the projection operator (Galkin et al., 2023) and its interplay with the multi-hop query answering framework. For instance, new mechanisms of tackling the multi-source propagation, better pre-training strategies, and scaling might positively impact the zero-shot CLQA performance.
@@
Conclusion and Future Work. We presented UltraQuery, the first approach for inductive zero-shot complex logical query answering on any KG that combines a parameterized, inductive projection operator with non-parametric logical operators. Alleviating the multi-source message propagation issue is the key to adapt pre-trained projection operators into the multi-hop query answering framework. UltraQuery performs comparably to or better than strong baselines trained specifically on each graph and at the same time retains key qualitative features like faithfullness and answer cardinality estimation. Having a single query answering model working on any KG, the scope for future work is vast as highlighted by Ren et al. (2023) and includes, for example, better theoretical understanding of logical expressiveness bounds, supporting more query patterns beyond simple trees (Yin et al., 2023a, b), queries without anchor nodes (Barceló et al., 2023), hyper-relational queries (Alivanistos et al., 2022), queries with numerical literals (Demir et al., 2023), or temporal queries (Lin et al., 2023).
Impact Statement
@@
We do not envision direct ethical or societal consequences of this work. Still, models capable of zero-shot inference on any graph might be applied to domains other than those designed by the authors. Positive impacts include saving compute resources and reducing carbon footprint of training specific models tailored for each graph.
Acknowledgments
JZ and BR are supported in part by the National Science Foundation (NSF) awards, CCF-1918483, CAREER IIS-1943364 and CNS-2212160, Amazon Research Award, AnalytiXIN, and the Wabash Heartland Innovation Network (WHIN). Computing infrastructure was supported in part by CNS-1925001 (CloudBank). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Acknowledgments
JT and ZZ are supported by Intel-Mila partnership program, the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund and a NRC Collaborative R&D Project (AI4D-CORE-06). This project was also partially funded by IVADO Fundamental Research Project grant PRF-2019-3583139727. The computation resource of this project is supported by Mila [^3] https://mila.quebec/, Calcul Québec [^4] https://www.calculquebec.ca/ and the Digital Research Alliance of Canada [^5] https://alliancecan.ca/.
@@ A.1 Dataset statistics
First, Table 4, Table 4, and Table 4 provide the necessary details on the graphs behind the CLQA datasets. Then, Table 7, Table 7, and Table 7 list the query statistics. Transductive datasets are BetaE (Ren & Leskovec, 2020) datasets, inductive $(e)$ datasets are adopted from Galkin et al. (2022b) where validation and test inference graphs extend the training graph. The ratio denotes the size of the inference graph to the size of the training graph (in the number of nodes), that is, ${\mathcal{V}}_{\textit{inf}}/{\mathcal{V}}_{\textit{train}}$. In the following Section A.2 we provide more details in sampling 11 new inductive $(e,r)$ datasets WikiTopics-CLQA.
Table 2: Graph in transductive datasets (3) from Ren & Leskovec (2020). Inverse triples and edge types are included in the splits. Train, Valid, Test denote triples in the respective set.
@@
Table 3: Graphs in inductive $(e)$ datasets (9) from Galkin et al. (2022b). Inverse triples and edge types are included in all graphs. Validation and Test splits contain an inference graph $({\mathcal{V}}_{\textit{inf}},{\mathcal{E}}_{\textit{inf}})$ which is a superset of the training graph with new nodes, and missing edges to predict (Valid and Test, respectively).
Ratio, % Rels Training Graph Validation Graph Test Graph
Entities Triples Entities Triples Valid Entities Triples Test
106% 466 13,091 493,425 13,801 551,336 10,219 13,802 538,896 8,023
113% 468 11,601 401,677 13,022 491,518 15,849 13,021 486,068 14,893
122% 466 10,184 298,879 12,314 413,554 20,231 12,314 430,892 23,289
134% 466 8,634 228,729 11,468 373,262 25,477 11,471 367,810 24,529
150% 462 7,232 162,683 10,783 311,462 26,235 10,782 331,352 29,755
175% 436 5,560 102,521 9,801 265,412 28,691 9,781 266,494 28,891
217% 446 4,134 52,455 9,062 227,284 30,809 9,058 212,386 28,177
300% 412 2,650 24,439 8,252 178,680 27,135 8,266 187,156 28,657
550% 312 1,084 5,265 7,247 136,558 22,981 7,275 133,524 22,503
@@
Dataset Training Graph Validation Graph Test Graph
Entities Rels Triples Entities Rels Triples Valid Entities Rels Triples Test
Art 10000 65 27262 10000 65 27262 3026 10000 65 28023 3113
Award 10000 17 23821 10000 13 23821 2646 10000 17 25056 2783
Education 10000 19 14355 10000 19 14355 1594 10000 19 14193 1575
Health 10000 31 15539 10000 31 15539 1725 10000 31 15337 1703
Infrastructure 10000 37 21990 10000 37 21990 2443 10000 37 21646 2405
Location 10000 62 85063 10000 62 85063 9451 10000 62 80269 8917
Organization 10000 34 33325 10000 34 33325 3702 10000 34 31314 3357
People 10000 40 55698 10000 40 55698 6188 10000 40 58530 6503
Science 10000 66 12576 10000 66 12576 1397 10000 66 12516 1388
Sport 10000 34 47251 10000 34 47251 5250 10000 34 46717 5190
Taxonomy 10000 59 18921 10000 59 18921 2102 10000 59 19416 2157
@@
Table 7: WikiTopics-CLQA statistics: the number of queries generated per query pattern for each topic knowledge graph of WikiTopics (Gao et al., 2023). Numbers are the same for both the training and inference graph. We follow the same 14 query patterns introduced by Ren & Leskovec (2020).
@@ A.2 WikiTopics-CLQA
The WikiTopics dataset introduced by Gao et al. (2023) was used to evaluate link prediction model’s zero-shot performance in the inductive $(e,r)$ setting, i.e., when the test-time inference graph contains both new entities and new relations unseen in training. It grouped relations into 11 different topics, or domains, such as art, education, health care, and sport. Two graphs, $\mathcal{G}_{\textit{train}}^{(T)}$ and $\mathcal{G}_{\textit{inf}}^{(T)}$ along with the missing triples $\mathcal{E}^{(T)}_{\textit{valid}}$ and $\mathcal{E}^{(T)}_{\textit{test}}$, were provided for each topic $T$, which had the same set of relations but different (potentially overlapping) set of entities. The goal was to train models on the training graphs $\mathcal{G}_{\textit{train}}^{(T)}$ of some topic $T$, and test on the inference graph $\mathcal{G}_{\textit{inf}}^{(T^{\prime})}$ of an unseen topic $T^{\prime}$. The model’s validation performance was evaluated on the missing triples $\mathcal{E}^{(T)}_{\textit{valid}}$ when observing training graph $\mathcal{G}_{\textit{train}}^{(T)}$ as inputs, and its test performance was evaluated on $\mathcal{E}^{(T^{\prime})}_{\textit{test}}$ when observing the test inference graph $\mathcal{G}_{\textit{inf}}^{(T^{\prime})}$ as inputs. Table 4 shows the statistics of the 11 topic-specific knowledge graphs in WikiTopics.
@@ Appendix B Hyperparameters and Baselines
Both UltraQuery and UltraQuery LP are initialized from the pre-trained Ultra checkpoint published by the authors and have the same GNN architectures with parameter count (177k). UltraQuery is further trained for 10,000 steps on complex queries from the FB15k237 dataset. Hyperparameters of UltraQuery and its training details are listed in Table 8.
@@
Transductive (3 datasets): QTO (Bai et al., 2023). QTO requires 2000dd ComplEx (Trouillon et al., 2016) entity and relation embeddings pre-computed for each graph, e.g., taking 30M parameters on the FB15k237 graph with 15k nodes. Further, QTO materializes the whole $({\mathcal{V}}\times{\mathcal{V}}\times{\mathcal{R}})$ 3D tensor of scores of all possible triples for each graph. Pre-computing such tensors on three datasets takes considerable space and time, e.g., 8 hours for FB15k with heavy sparsification settings to fit onto a 24 GB GPU.
Inductive $(e)$ (9 datasets): GNN-QE (Zhu et al., 2022). The framework of GNN-QE is similar to UltraQuery but the backbone relation projection operator is implemented with NBFNet (Zhu et al., 2021) which is only inductive to entities and still learns graph-specific relation embeddings. Parameter count, therefore, depends on the number of unique relation types, e.g., 2M for the FB 175% split with 436 unique relations.
Inductive $(e,r)$ (11 datasets): for newly sampled datasets, due to the absence of fully inductive trainable baselines, we compare against a rule-based heuristic baseline similar to the baseline in Galkin et al. (2022b). The baseline looks up candidate entities that have the same incoming relation type as the current relation projection (note that the identity of the starting head node in this case is ignored). The heuristic filters entities into two classes (satisfying the incoming relation type or not), hence, in order to create a ranking, we randomly shuffle entities in each class. This baseline is non-parametric, does not require any training, and represents a sanity check of CLQA models. Still, as shown in Galkin et al. (2022b), the baseline might outperform certain inductive reasoning approaches parameterized by neural networks.
@@ Appendix C More Results
UltraQuery LP UltraQuery Best Baseline EPFO avg Negation avg EPFO avg Negation avg EPFO avg Negation avg MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 transductive datasets FB15k237 0.221 0.379 0.081 0.165 0.252 0.395 0.090 0.195 0.335 0.491 0.155 0.295 FB15k 0.523 0.678 0.341 0.507 0.525 0.668 0.239 0.437 0.740 0.837 0.492 0.664 NELL995 0.256 0.405 0.070 0.139 0.228 0.339 0.067 0.147 0.329 0.483 0.129 0.268 inductive $(e)$ datasets FB 106% 0.204 0.353 0.056 0.119 0.244 0.396 0.110 0.210 0.373 0.504 0.256 0.377 FB 113% 0.226 0.377 0.069 0.148 0.258 0.407 0.123 0.228 0.352 0.494 0.214 0.339 FB 121% 0.253 0.418 0.083 0.171 0.296 0.456 0.144 0.283 0.323 0.462 0.173 0.291 FB 133% 0.275 0.442 0.086 0.170 0.308 0.464 0.135 0.263 0.353 0.514 0.197 0.341 FB 150% 0.283 0.456 0.094 0.195 0.323 0.482 0.150 0.288 0.339 0.493 0.167 0.303 FB 175% 0.309 0.488 0.107 0.215 0.346 0.510 0.175 0.322 0.351 0.507 0.188 0.336 FB 217% 0.338 0.516 0.121 0.237 0.373 0.536 0.194 0.349 0.346 0.492 0.174 0.301 FB 300% 0.328 0.513 0.109 0.220 0.365 0.528 0.184 0.340 0.291 0.426 0.125 0.224 FB 550% 0.346 0.527 0.122 0.236 0.375 0.536 0.193 0.337 0.222 0.331 0.091 0.158 inductive $(e,r)$ datasets Art 0.207 0.342 0.055 0.112 0.237 0.333 0.087 0.136 0.016 0.031 0.006 0.014 Award 0.192 0.353 0.018 0.034 0.178 0.287 0.108 0.193 0.004 0.006 0.002 0.002 Edu 0.137 0.222 0.052 0.104 0.187 0.257 0.094 0.150 0.008 0.014 0.003 0.005 Health 0.316 0.420 0.131 0.187 0.298 0.382 0.265 0.360 0.019 0.040 0.010 0.020 Infrastructure 0.362 0.516 0.095 0.195 0.370 0.455 0.175 0.244 0.001 0.018 0.003 0.005 Location 0.388 0.540 0.178 0.306 0.434 0.538 0.240 0.374 0.007 0.017 0.001 0.002 Organization 0.137 0.259 0.002 0.042 0.215 0.355 0.120 0.253 0.005 0.008 0.001 0.002 People 0.361 0.498 0.160 0.262 0.377 0.465 0.213 0.266 0.005 0.010 0.002 0.003 Science 0.207 0.342 0.045 0.092 0.236 0.356 0.093 0.172 0.041 0.085 0.010 0.020 Sport 0.148 0.259 0.049 0.092 0.202 0.335 0.145 0.229 0.015 0.030 0.005 0.008 Taxonomy 0.178 0.256 0.119 0.196 0.275 0.348 0.211 0.306 0.034 0.066 0.009 0.017

### 01 Oct 2024 v2

@@ 6 Discussion and Future Work
Limitations. The only parameterized component of UltraQuery is the projection operator and, therefore, limitations and improvement opportunities stem from the projection operator [15] and its interplay with the multi-hop query answering framework. For instance, new mechanisms of tackling the multi-source propagation, better pre-training strategies, and scaling might positively impact the zero-shot CLQA performance. The support for very large KGs could be further improved by adopting more scalable entity-level GNN predictors like A*Net [44] or AdaProp [39] which have been shown to scale to graphs of millions of nodes. We are optimistic that UltraQuery could scale to such graphs when integrated with those models.
@@
Conclusion and Future Work. We presented UltraQuery, the first foundation model for inductive zero-shot complex logical query answering on any KG that combines a parameterized, inductive projection operator with non-parametric logical operators. Alleviating the multi-source message propagation issue is the key to adapt pre-trained projection operators into the multi-hop query answering framework. UltraQuery performs comparably to or better than strong baselines trained specifically on each graph and at the same time retains key qualitative features like faithfullness and answer cardinality estimation. Having a single query answering model working on any KG, the scope for future work is vast as highlighted by Ren et al. [27] and includes, for example, better theoretical understanding of logical expressiveness bounds, supporting more query patterns beyond simple trees [35, 36], queries without anchor nodes [7], hyper-relational queries [1], queries with numerical literals [11], or temporal queries [23].
@@
Impact Statement. We do not envision direct ethical or societal consequences of this work. Still, models capable of zero-shot inference on any graph might be applied to domains other than those designed by the authors. Positive impacts include saving compute resources and reducing carbon footprint of training specific models tailored for each graph.
Acknowledgements
@@
This project is supported by Intel-Mila partnership program, the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund and a NRC Collaborative R&D Project (AI4D-CORE-06). This project was also partially funded by IVADO Fundamental Research Project grant PRF-2019-3583139727. The computation resource of this project is supported by Mila [^3] https://mila.quebec/, Calcul Québec [^4] https://www.calculquebec.ca/ and the Digital Research Alliance of Canada [^5] https://alliancecan.ca/.
This work was funded in part by the National Science Foundation (NSF) awards, CCF-1918483, CAREER IIS-1943364 and CNS-2212160, Amazon Research Award, AnalytiXIN, and the Wabash Heartland Innovation Network (WHIN). Computing infrastructure was supported in part by CNS-1925001 (CloudBank). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ A.1 Dataset statistics
First, Table 6, Table 6, and Table 6 provide the necessary details on the graphs behind the CLQA datasets. Then, Table 9, Table 9, and Table 9 list the query statistics. Transductive datasets are BetaE [25] datasets (MIT license), inductive $(e)$ datasets are adopted from Galkin et al. [14] (CC BY 4.0 license) where validation and test inference graphs extend the training graph. The ratio denotes the size of the inference graph to the size of the training graph (in the number of nodes), that is, ${\mathcal{V}}_{\textit{inf}}/{\mathcal{V}}_{\textit{train}}$. In the following Section A.2 we provide more details in sampling 11 new inductive $(e,r)$ datasets WikiTopics-CLQA (available under the CC BY 4.0 license).
Table 4: Graph in transductive datasets (3) from Ren and Leskovec [25]. Inverse triples and edge types are included in the splits. Train, Valid, Test denote triples in the respective set.
@@
Table 5: Graphs in inductive $(e)$ datasets (9) from Galkin et al. [14]. Inverse triples and edge types are included in all graphs. Validation and Test splits contain an inference graph $({\mathcal{V}}_{\textit{inf}},{\mathcal{E}}_{\textit{inf}})$ which is a superset of the training graph with new nodes, and missing edges to predict (Valid and Test, respectively).
Ratio, % Rels Training Graph Validation Graph Test Graph Entities Triples Entities Triples Valid Entities Triples Test 106% 466 13,091 493,425 13,801 551,336 10,219 13,802 538,896 8,023 113% 468 11,601 401,677 13,022 491,518 15,849 13,021 486,068 14,893 122% 466 10,184 298,879 12,314 413,554 20,231 12,314 430,892 23,289 134% 466 8,634 228,729 11,468 373,262 25,477 11,471 367,810 24,529 150% 462 7,232 162,683 10,783 311,462 26,235 10,782 331,352 29,755 175% 436 5,560 102,521 9,801 265,412 28,691 9,781 266,494 28,891 217% 446 4,134 52,455 9,062 227,284 30,809 9,058 212,386 28,177 300% 412 2,650 24,439 8,252 178,680 27,135 8,266 187,156 28,657 550% 312 1,084 5,265 7,247 136,558 22,981 7,275 133,524 22,503
@@
Dataset Training Graph Validation Graph Test Graph Entities Rels Triples Entities Rels Triples Valid Entities Rels Triples Test Art 10000 65 27262 10000 65 27262 3026 10000 65 28023 3113 Award 10000 17 23821 10000 13 23821 2646 10000 17 25056 2783 Education 10000 19 14355 10000 19 14355 1594 10000 19 14193 1575 Health 10000 31 15539 10000 31 15539 1725 10000 31 15337 1703 Infrastructure 10000 37 21990 10000 37 21990 2443 10000 37 21646 2405 Location 10000 62 85063 10000 62 85063 9451 10000 62 80269 8917 Organization 10000 34 33325 10000 34 33325 3702 10000 34 31314 3357 People 10000 40 55698 10000 40 55698 6188 10000 40 58530 6503 Science 10000 66 12576 10000 66 12576 1397 10000 66 12516 1388 Sport 10000 34 47251 10000 34 47251 5250 10000 34 46717 5190 Taxonomy 10000 59 18921 10000 59 18921 2102 10000 59 19416 2157
@@
Table 9: WikiTopics-CLQA statistics: the number of queries generated per query pattern for each topic knowledge graph of WikiTopics [16]. Numbers are the same for both the training and inference graph. We follow the same 14 query patterns introduced by Ren and Leskovec [25].
@@ A.2 WikiTopics-CLQA
The WikiTopics dataset introduced by Gao et al. [16] was used to evaluate link prediction model’s zero-shot performance in the inductive $(e,r)$ setting, i.e., when the test-time inference graph contains both new entities and new relations unseen in training. It grouped relations into 11 different topics, or domains, such as art, education, health care, and sport. Two graphs, $\mathcal{G}_{\textit{train}}^{(T)}$ and $\mathcal{G}_{\textit{inf}}^{(T)}$ along with the missing triples $\mathcal{E}^{(T)}_{\textit{valid}}$ and $\mathcal{E}^{(T)}_{\textit{test}}$, were provided for each topic $T$, which had the same set of relations but different (potentially overlapping) set of entities. The goal was to train models on the training graphs $\mathcal{G}_{\textit{train}}^{(T)}$ of some topic $T$, and test on the inference graph $\mathcal{G}_{\textit{inf}}^{(T^{\prime})}$ of an unseen topic $T^{\prime}$. The model’s validation performance was evaluated on the missing triples $\mathcal{E}^{(T)}_{\textit{valid}}$ when observing training graph $\mathcal{G}_{\textit{train}}^{(T)}$ as inputs, and its test performance was evaluated on $\mathcal{E}^{(T^{\prime})}_{\textit{test}}$ when observing the test inference graph $\mathcal{G}_{\textit{inf}}^{(T^{\prime})}$ as inputs. Table 6 shows the statistics of the 11 topic-specific knowledge graphs in WikiTopics.
@@ Appendix B Hyperparameters and Baselines
Both UltraQuery and UltraQuery LP are implemented with PyTorch [24] (BSD-style license) and PyTorch-Geometric [12] (MIT license). Both UltraQuery and UltraQuery LP are initialized from the pre-trained Ultra checkpoint published by the authors and have the same GNN architectures with parameter count (177k). UltraQuery is further trained for 10,000 steps on complex queries from the FB15k237 dataset. Hyperparameters of UltraQuery and its training details are listed in Table 10.
@@
Transductive (3 datasets): QTO [5]. QTO requires 2000dd ComplEx [30] entity and relation embeddings pre-computed for each graph, e.g., taking 30M parameters on the FB15k237 graph with 15k nodes. Further, QTO materializes the whole $({\mathcal{V}}\times{\mathcal{V}}\times{\mathcal{R}})$ 3D tensor of scores of all possible triples for each graph. Pre-computing such tensors on three datasets takes considerable space and time, e.g., 8 hours for FB15k with heavy sparsification settings to fit onto a 24 GB GPU.
Inductive $(e)$ (9 datasets): GNN-QE [43]. The framework of GNN-QE is similar to UltraQuery but the backbone relation projection operator is implemented with NBFNet [42] which is only inductive to entities and still learns graph-specific relation embeddings. Parameter count, therefore, depends on the number of unique relation types, e.g., 2M for the FB 175% split with 436 unique relations.
Inductive $(e,r)$ (11 datasets): for newly sampled datasets, due to the absence of fully inductive trainable baselines, we compare against a rule-based heuristic baseline similar to the baseline in Galkin et al. [14]. The baseline looks up candidate entities that have the same incoming relation type as the current relation projection (note that the identity of the starting head node in this case is ignored). The heuristic filters entities into two classes (satisfying the incoming relation type or not), hence, in order to create a ranking, we randomly shuffle entities in each class. This baseline is non-parametric, does not require any training, and represents a sanity check of CLQA models. Still, as shown in Galkin et al. [14], the baseline might outperform certain inductive reasoning approaches parameterized by neural networks.
@@ Appendix C More Results
UltraQuery LP UltraQuery Best Baseline EPFO avg Negation avg EPFO avg Negation avg EPFO avg Negation avg MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 transductive datasets FB15k237 0.216 0.362 0.082 0.164 0.242 0.378 0.08 0.174 0.335 0.491 0.155 0.295 FB15k 0.501 0.672 0.291 0.465 0.764 0.834 0.567 0.725 0.74 0.837 0.492 0.664 NELL995 0.249 0.395 0.079 0.16 0.226 0.341 0.073 0.159 0.329 0.483 0.129 0.268 inductive $(e)$ datasets FB 550% 0.340 0.518 0.134 0.251 0.373 0.535 0.186 0.332 0.222 0.331 0.091 0.158 FB 300% 0.320 0.496 0.117 0.227 0.359 0.526 0.168 0.312 0.291 0.426 0.125 0.224 FB 217% 0.332 0.509 0.133 0.252 0.375 0.537 0.186 0.337 0.346 0.492 0.174 0.301 FB 175% 0.297 0.469 0.110 0.214 0.338 0.499 0.159 0.297 0.351 0.507 0.188 0.336 FB 150% 0.279 0.445 0.097 0.190 0.316 0.473 0.132 0.255 0.339 0.493 0.167 0.303 FB 133% 0.266 0.426 0.087 0.167 0.298 0.451 0.122 0.238 0.353 0.514 0.197 0.341 FB 121% 0.246 0.400 0.081 0.164 0.279 0.430 0.119 0.232 0.323 0.462 0.173 0.291 FB 113% 0.217 0.362 0.067 0.136 0.240 0.387 0.097 0.192 0.352 0.494 0.214 0.339 FB 106% 0.200 0.340 0.054 0.114 0.226 0.370 0.086 0.162 0.373 0.504 0.256 0.377 inductive $(e,r)$ datasets Art 0.249 0.389 0.086 0.157 0.248 0.349 0.083 0.137 0.016 0.031 0.006 0.014 Award 0.224 0.413 0.046 0.098 0.227 0.354 0.152 0.274 0.004 0.006 0.002 0.002 Edu 0.142 0.258 0.066 0.122 0.179 0.249 0.119 0.176 0.008 0.014 0.003 0.005 Health 0.317 0.466 0.159 0.231 0.317 0.394 0.419 0.525 0.019 0.040 0.010 0.020 Infrastructure 0.392 0.551 0.087 0.170 0.337 0.461 0.235 0.356 0.010 0.018 0.003 0.005 Location 0.508 0.678 0.198 0.371 0.569 0.679 0.402 0.585 0.007 0.017 0.001 0.002 Organization 0.098 0.190 0.023 0.048 0.169 0.270 0.082 0.171 0.005 0.008 0.001 0.002 People 0.373 0.530 0.182 0.281 0.332 0.443 0.194 0.285 0.005 0.010 0.002 0.003 Science 0.206 0.348 0.048 0.093 0.158 0.255 0.071 0.131 0.041 0.085 0.010 0.020 Sport 0.215 0.357 0.095 0.166 0.252 0.371 0.178 0.269 0.015 0.030 0.005 0.008 Taxonomy 0.230 0.315 0.151 0.255 0.290 0.360 0.183 0.261 0.034 0.066 0.009 0.017